# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ 6.2.2. Recovery
Many studies have explored recovery techniques to improve UAV system availability and minimize downtime after faults. Lightweight task migration and redundancy-based switching strategies are frequently employed to reallocate computing workloads in response to partial system failures or degraded node performance (Luo et al., 2022; Zuo et al., 2022; Luo et al., 2021b; Awada et al., 2023b; Pan et al., 2022). In scenarios where UAVs operate in harsh or disconnected environments, distributed task re-execution or fallback mechanisms are introduced to ensure mission continuity (Callegaro et al., 2021; Choutri et al., 2022; Hou et al., 2020d; Machida et al., 2023a, b). Energy-aware recovery approaches are also proposed, balancing the trade-off between recovery success and remaining UAV resources (Wang et al., 2024b; Yu et al., 2022b; Zhang et al., 2024b; Watanabe and Machida, 2022).
@@ 6.2.3. Redundant Architecture
Redundant architecture refers to fault tolerance strategies where multiple instances of hardware, software components, or communication paths are integrated into UAV-based systems to maintain service continuity in the presence of faults. Rather than preventing faults from occurring, these techniques aim to ensure that system functionality is preserved despite failures in some components. In UAV-based systems, redundancy is employed across communication links, computation nodes, sensors, and actuator subsystems. For example, dual communication modules and duplicated computing resources are used to tolerate communication failures and processing overloads without interrupting mission-critical services (Ma et al., 2024; Farrukh and West, 2023; Kilbourne et al., 2021; Awada et al., 2021b; Karam et al., 2024a).
Multi-UAV swarms with overlapping coverage enable persistent sensing and data transmission even when individual units become inoperative (Patnayak et al., 2021a; Zhang et al., 2021c; Bian et al., 2023; Jung et al., 2023b; Kim, 2024). Backup control modules and redundant edge/fog nodes are also integrated to support recovery and seamless task offloading in the event of partial system degradation or network instability (Kaleem et al., 2019b; Dogea et al., 2023; Callegaro et al., 2021; Pacheco et al., 2021b; Petrenko et al., 2023; El Sayed et al., 2023; Sara et al., 2016). Additionally, hybrid cloud-edge frameworks offer parallel computational paths to guarantee task execution continuity under localized failures or load fluctuations (Luo et al., 2020; La Salandra et al., 2024; Khan et al., 2022; Chen and Ma, 2024; Wang et al., 2020a; Shen et al., 2023a; Tang et al., 2023). These fault-tolerant architectures directly contribute to dependability attributes such as availability, reliability, and resilience by enabling uninterrupted operations and graceful degradation in the presence of faults, rather than complete system failure.
@@ 6.2.4. Network Topology Design
Network Topology Design refers to fault tolerance strategies that focus on maintaining reliable and resilient communication within systems, especially in environments where signal disruption, disconnection, or routing instability is common. Rather than preventing communication faults outright, these techniques aim to sustain service availability and network functionality even when individual links or nodes fail during operation. In UAV-based systems, dynamic topologies such as flying ad hoc networks (FANETs), mesh structures, and mobile relays are often employed to mitigate link degradation and maintain multi-hop connectivity. For instance, adaptive topology control and link prediction mechanisms are used to anticipate and respond to potential disconnections due to UAV mobility or energy depletion (Xu et al., 2021; Guruprasad et al., 2024; Zhang et al., 2023b; Vamvakas et al., 2019; Chen et al., 2022c; Wang et al., 2021e; Dao et al., 2021; Ozger et al., 2018b; Su, 2021; Wang et al., 2022c).
UAV-aided reconfiguration is another popular approach, where drones are repositioned to act as temporary relays or mobile base stations to bridge coverage gaps and reduce the risk of total communication failure (Zhou et al., 2024b; Kong et al., 2022a; Zhang et al., 2021a; Mukherjee et al., 2020; Nwadiugwu et al., 2021; Mukherjee et al., 2023; Yue et al., 2018; Wang et al., 2020b; Zhou et al., 2024c; Wang et al., 2019b). These topology-aware fault tolerance mechanisms directly contribute to availability, connectivity, and system resilience, ensuring that UAV missions such as remote surveillance, disaster response, or public safety operations can continue despite disruptions in the communication infrastructure.
@@ 6.3. Fault Removal
Diagnosis | (Xu and Zhai, 2020), (Venkatagiri et al., 2018b), (Deng et al., 2021), (Jo et al., 2018), (Amin and Obermaisser, 2018a), (Ramasamy et al., 2024), (Zhang and Zhang, 2024), (Wang et al., 2021d), (Wang et al., 2022a), (Wang et al., 2024d), (Kostoglotov et al., 2024), (Bobronnikov et al., 2022), (Bouzidi et al., 2020), (Amin and Obermaisser, 2018b) | |
Correction | (Gu et al., 2022b), (Luo et al., 2022), (Bobronnikov et al., 2022) | |
@@ 6.3.2. Diagnosis:
Diagnosis primarily addresses software faults, hardware faults, and network disconnection, where faults may manifest as unexpected delays, abnormal UAV motion, or degraded communication quality. Several studies utilize behavioral and model-based diagnosis to identify operator-induced errors by comparing them with expected control models during simulated missions, providing insights into human-machine interaction errors and their safety implications (Xu and Zhai, 2020; Amin and Obermaisser, 2018a). Secondly, cloud and edge-enabled anomaly detection frameworks are increasingly employed to enhance scalability and responsiveness, particularly for large-scale UAV applications such as power line inspection and predictive maintenance. These include distributed isolation forest algorithms and embedded monitoring tools that enable timely fault isolation while minimizing computational burden (Ramasamy et al., 2024; Zhang and Zhang, 2024; Wang et al., 2022a).
On the other hand, intelligent agents and software-based diagnostic systems support modular and autonomous operation by integrating distributed reasoning or software approximation techniques. For example, agent-based building inspection platforms and approximation-aware video summarization systems leverage real-time assessment capabilities to maintain operational resilience and energy efficiency (Venkatagiri et al., 2018b; Jo et al., 2018). Meanwhile, cooperative and network-level diagnosis approaches utilize drone-assisted sensing and edge intelligence to detect anomalies within broader communication networks, including space-air-ground infrastructures, thereby ensuring systemic fault awareness and coordination (Deng et al., 2021; Wang et al., 2021d).
@@ 6.4. Fault Forecasting
Fault Prevention | Ordinal Evaluation | (Bouzidi et al., 2020), (Guo et al., 2022b), (Hossain et al., 2017), (Manzini et al., 2023b), (Fesenko et al., 2023), (Venkatagiri et al., 2018b), (Kramer and Stubberud, 2021), (Cao et al., 2021) |
Probabilistic Evaluation | (Sabino et al., 2024), (Ngoenriang et al., 2022b). (Goel et al., 2021), (Tahat et al., 2020), (Ji et al., 2024), (Apostolopoulos et al., 2023), (Bai et al., 2023b), (Brito et al., 2021), (Falcão et al., 2024), (Machida and Andrade, 2021a), (Kim and Kim, 2021), (Manzoor et al., 2021), (Thapliyal et al., 2024a), (Machida et al., 2023a), (Dai et al., 2018), (She and Ouyang, 2021), (Machida et al., 2023b), (Chen et al., 2021b), (Yang et al., 2017), (Falcão et al., 2023a), (Watanabe and Machida, 2022), (Schäfer et al., 2018), (Jeong et al., 2020a), (Picano and Fantacci, 2024a), (Kabashkin, 2023), (Yao and Ansari, 2018), (She et al., 2018), (Silva et al., 2024), (Hu and Seiler, 2015), (Falcão et al., 2023b), (Li et al., 2019), (Bai et al., 2023a), (Venkatagiri et al., 2018a), (Zhang, 2020) |
@@ 6.4.2. Probabilistic Evaluation:
In the context of dependability of UAV-based systems, probabilistic evaluation directly supports reliability analysis, availability estimation, and maintenance planning, offering a mathematical foundation for system robustness under varying workloads and fault scenarios. For example, studies that apply probabilistic models to assess component failure rates and communication link interruptions under uncertain conditions (Sabino et al., 2024; Ngoenriang et al., 2022b; Tahat et al., 2020). Other studies integrate Markov models and stochastic processes to estimate system uptime and mission success rates (Ji et al., 2024; Apostolopoulos et al., 2023; Bai et al., 2023b). Bayesian networks and Monte Carlo simulations are used to analyze the effects of uncertain environmental conditions or resource variability on UAV operations (Falcão et al., 2024; Kim and Kim, 2021; Manzoor et al., 2021).
@@ 7.2. Constellation of satellites
To address these, reliability considerations involve energy-efficient designs, such as low-power transceivers and burst transmissions, as well as hybrid communication strategies that combine satellite and terrestrial links (Wang et al., 2019d) with fallback mechanisms. Robust signal resilience could be realised through adaptive modulation, latency-aware decision-making, and security measures. While satellite-enabled UAVs offer unprecedented operational range and flexibility, their deployment requires careful architectural design and robust fault-tolerance mechanisms to maintain dependable performance in dynamic and contested environments.
@@ 7.3. B5G/6G network integration
Future research should explore the design of UAV-based systems that are tailored to B5G/6G environments (Shrestha et al., 2021). In addition, joint optimization of communication and computing, including seamless handovers and UAV trajectory planning with connectivity guarantees, will be crucial to maintain QoS under highly dynamic aerial conditions. Integration with terahertz communication, intelligent reflecting surfaces (IRS), and network function virtualization (NFV) can further extend the coverage, reduce energy consumption, and improve resilience of UAV networks (Chen et al., 2021c).
@@ Acknowledgements.
- Akbari et al. (2024a) Mohammad Akbari, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2024a. AoI-aware energy-efficient SFC in UAV-aided smart agriculture using asynchronous federated learning. IEEE Open J. Commun. Soc. 5 (2024), 1222–1242.
- Akbari et al. (2024b) Mohammad Akbari, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2024b. AoI-aware energy-efficient SFC in UAV-aided smart agriculture using asynchronous federated learning. IEEE Open J. Commun. Soc. 5 (2024), 1222–1242.
@@
- Amin and Obermaisser (2018a) Sarah Amin and Roman Obermaisser. 2018a. A time-triggered scheduling algorithm for active diagnosis in heterogeneous distributed systems. In 2018 IEEE International Conference on Computational Science and Engineering (CSE) (Bucharest). IEEE.
- Amin and Obermaisser (2018b) Sarah Amin and Roman Obermaisser. 2018b. A time-triggered scheduling algorithm for active diagnosis in heterogeneous distributed systems. In 2018 IEEE International Conference on Computational Science and Engineering (CSE) (Bucharest). IEEE.
@@
- Awada et al. (2021a) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2021a. Air-to-air collaborative learning: A multi-task orchestration in federated aerial computing. In 2021 IEEE 14th International Conference on Cloud Computing (CLOUD) (Chicago, IL, USA). IEEE.
- Awada et al. (2021b) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2021b. Air-to-air collaborative learning: A multi-task orchestration in federated aerial computing. In 2021 IEEE 14th International Conference on Cloud Computing (CLOUD) (Chicago, IL, USA). IEEE.
- Awada et al. (2022a) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2022a. AirEdge: A dependency-aware multi-task orchestration in federated aerial computing. IEEE Trans. Veh. Technol. 71, 1 (Jan. 2022), 805–819.
- Awada et al. (2022b) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2022b. AirEdge: A dependency-aware multi-task orchestration in federated aerial computing. IEEE Trans. Veh. Technol. 71, 1 (Jan. 2022), 805–819.
- Awada et al. (2023a) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, Shuangzhi Li, and Shouyi Yang. 2023a. EdgeDrones: Co-scheduling of drones for multi-location aerial computing missions. J. Netw. Comput. Appl. 215, 103632 (June 2023), 103632.
- Awada et al. (2023b) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, Shuangzhi Li, and Shouyi Yang. 2023b. EdgeDrones: Co-scheduling of drones for multi-location aerial computing missions. J. Netw. Comput. Appl. 215, 103632 (June 2023), 103632.
@@
- Bai et al. (2023c) Haohao Bai, Zhineng Luo, Meihui Chen, and Senke Gui. 2023c. Energy saving research on auxiliary task unloading and resource allocation of UAV based on edge computing. In 2023 International Conference on Communications, Computing and Artificial Intelligence (CCCAI) (Shanghai, China). IEEE, 103–108.
@@
- Bai et al. (2023b) Jing Bai, Xiaolin Chang, Ricardo J Rodríguez, Kishor S Trivedi, and Shupan Li. 2023b. Towards UAV-based MEC service chain resilience evaluation: A quantitative modeling approach. IEEE Trans. Veh. Technol. 72, 4 (April 2023), 5181–5194.
- Bakhtiari et al. (2020a) Zohreh Bakhtiari, Rozita Jamili Oskouei, Mona Soleymani, and Akhtar Hussain Jalbani. 2020a. Presenting an effective method to detect and track the broken path in VANET using UAVs. Wirel. Commun. Mob. Comput. 2020 (Aug. 2020), 1–12.
- Bakhtiari et al. (2020b) Zohreh Bakhtiari, Rozita Jamili Oskouei, Mona Soleymani, and Akhtar Hussain Jalbani. 2020b. Presenting an effective method to detect and track the broken path in VANET using UAVs. Wirel. Commun. Mob. Comput. 2020 (Aug. 2020), 1–12.
@@
- Chen et al. (2022b) Haowei Chen, Liekang Zeng, Xiaoxi Zhang, and Xu Chen. 2022b. AdaDrone: Quality of navigation based neural adaptive scheduling for edge-assisted drones. In 2022 IEEE 42nd International Conference on Distributed Computing Systems (ICDCS) (Bologna, Italy). IEEE.
- Chen et al. (2022c) Lunyuan Chen, Rui Zhao, Ke He, Zichao Zhao, and Liseng Fan. 2022c. Intelligent ubiquitous computing for future UAV-enabled MEC network systems. Cluster Comput. 25, 4 (Aug. 2022), 2417–2427.
@@
- Chen et al. (2021b) Yali Chen, Bo Ai, Yong Niu, Hongliang Zhang, and Zhu Han. 2021b. Energy-constrained computation offloading in space-air-ground integrated networks using distributionally robust optimization. IEEE Trans. Veh. Technol. 70, 11 (Nov. 2021), 12113–12125.
@@
- Chen et al. (2021c) Zhi Chen, Boyu Ning, Chong Han, Zhongbao Tian, and Shaoqian Li. 2021c. Intelligent Reflecting Surface Assisted Terahertz Communications toward 6G. arXiv:2104.02897 [cs.IT] https://arxiv.org/abs/2104.02897
@@
- Chu et al. (2021a) Liju Chu, Xuejun Li, Jia Xu, Azadeh Ghari Neiat, and Xiao Liu. 2021a. A holistic service provision strategy for drone-as-a-service in MEC-based UAV delivery. In 2021 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE.
- Chu et al. (2021b) Liju Chu, Xuejun Li, Jia Xu, Azadeh Ghari Neiat, and Xiao Liu. 2021b. A holistic service provision strategy for drone-as-a-service in MEC-based UAV delivery. In 2021 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE.
@@
- Falcão et al. (2023a) Marcos Falcão, Caio Souza, Andson Balieiro, and Kelvin Dias. 2023a. Dynamic resource allocation for URLLC in UAV-enabled multi-access edge computing. In 2023 Joint European Conference on Networks and Communications & 6G Summit (EuCNC/6G Summit) (Gothenburg, Sweden). IEEE.
- Falcão et al. (2023b) Marcos Falcão, Caio Souza, Andson Balieiro, and Kelvin Dias. 2023b. Dynamic resource allocation for URLLC in UAV-enabled multi-access edge computing. In 2023 Joint European Conference on Networks and Communications & 6G Summit (EuCNC/6G Summit) (Gothenburg, Sweden). IEEE.
@@
- Faraci et al. (2023a) Giuseppe Faraci, Santi Agatino Rizzo, and Giovanni Schembra. 2023a. Green edge intelligence for smart management of a FANET in disaster-recovery scenarios. IEEE Trans. Veh. Technol. 72, 3 (March 2023), 3819–3831.
- Faraci et al. (2023b) Giuseppe Faraci, Santi Agatino Rizzo, and Giovanni Schembra. 2023b. Green edge intelligence for smart management of a FANET in disaster-recovery scenarios. IEEE Trans. Veh. Technol. 72, 3 (March 2023), 3819–3831.
@@
- Fatima et al. (2024a) Nida Fatima, Paresh Saxena, and Giovanni Giambene. 2024a. Deep reinforcement learning based computation offloading for xURLLC services with UAV-assisted IoT-based multi-access edge computing system. Wirel. Netw. 30, 9 (Dec. 2024), 7275–7291.
- Fatima et al. (2024b) Nida Fatima, Paresh Saxena, and Giovanni Giambene. 2024b. Deep reinforcement learning based computation offloading for xURLLC services with UAV-assisted IoT-based multi-access edge computing system. Wirel. Netw. 30, 9 (Dec. 2024), 7275–7291.
@@
- Gao et al. (2019a) Ang Gao, Yansu Hu, Wei Liang, Yizhi Lin, Lixin Li, and Xu Li. 2019a. A QoE-oriented scheduling scheme for energy-efficient computation offloading in UAV cloud system. IEEE Access 7 (2019), 68656–68668.
- Gao et al. (2019b) Ang Gao, Yansu Hu, Wei Liang, Yizhi Lin, Lixin Li, and Xu Li. 2019b. A QoE-oriented scheduling scheme for energy-efficient computation offloading in UAV cloud system. IEEE Access 7 (2019), 68656–68668.
- Gao et al. (2021a) Ang Gao, Qi Wang, Wei Liang, and Zhiguo Ding. 2021a. Game combined multi-agent reinforcement learning approach for UAV assisted offloading. IEEE Trans. Veh. Technol. 70, 12 (Dec. 2021), 12888–12901.
- Gao et al. (2021b) Ang Gao, Qi Wang, Wei Liang, and Zhiguo Ding. 2021b. Game combined multi-agent reinforcement learning approach for UAV assisted offloading. IEEE Trans. Veh. Technol. 70, 12 (Dec. 2021), 12888–12901.
@@
- Geng et al. (2024a) Liwei Geng, Hongbo Zhao, and Changming Zou. 2024a. Stackelberg game-based computation offloading and pricing in UAV assisted vehicular networks. IEEE Trans. Reliab. (2024), 1–15.
- Geng et al. (2024b) Liwei Geng, Hongbo Zhao, and Changming Zou. 2024b. Stackelberg game-based computation offloading and pricing in UAV assisted vehicular networks. IEEE Trans. Reliab. (2024), 1–15.
@@
- Gu et al. (2022a) Xiaohui Gu, Guoan Zhang, Mingxing Wang, Wei Duan, Miaowen Wen, and Pin-Han Ho. 2022a. UAV-aided energy-efficient edge computing networks: Security offloading optimization. IEEE Internet Things J. 9, 6 (March 2022), 4245–4258.
- Gu et al. (2022b) Xiaohui Gu, Guoan Zhang, Mingxing Wang, Wei Duan, Miaowen Wen, and Pin-Han Ho. 2022b. UAV-aided energy-efficient edge computing networks: Security offloading optimization. IEEE Internet Things J. 9, 6 (March 2022), 4245–4258.

### v2

@@ 6.2.2. Recovery
Many studies have explored recovery techniques to improve UAV system availability and minimize downtime after faults. Lightweight task migration and redundancy-based switching strategies are frequently employed to reallocate computing workloads in response to partial system failures or degraded node performance (Luo et al., 2022; Zuo et al., 2022; Luo et al., 2021; Awada et al., 2023; Pan et al., 2022). In scenarios where UAVs operate in harsh or disconnected environments, distributed task re-execution or fallback mechanisms are introduced to ensure mission continuity (Callegaro et al., 2021; Choutri et al., 2022; Hou et al., 2020b; Machida et al., 2023, 2023). Energy-aware recovery approaches are also proposed, balancing the trade-off between recovery success and remaining UAV resources (Wang et al., 2024b; Yu et al., 2022b; Zhang et al., 2024b; Watanabe and Machida, 2022).
@@ 6.2.3. Redundant Architecture
Redundant architecture refers to fault tolerance strategies where multiple instances of hardware, software components, or communication paths are integrated into UAV-based systems to maintain service continuity in the presence of faults. Rather than preventing faults from occurring, these techniques aim to ensure that system functionality is preserved despite failures in some components. In UAV-based systems, redundancy is employed across communication links, computation nodes, sensors, and actuator subsystems. For example, dual communication modules and duplicated computing resources are used to tolerate communication failures and processing overloads without interrupting mission-critical services (Ma et al., 2024; Farrukh and West, 2023; Kilbourne et al., 2021; Awada et al., 2021; Karam et al., 2024).
Multi-UAV swarms with overlapping coverage enable persistent sensing and data transmission even when individual units become inoperative (Patnayak et al., 2021; Zhang et al., 2021c; Bian et al., 2023; Jung et al., 2023a; Kim, 2024). Backup control modules and redundant edge/fog nodes are also integrated to support recovery and seamless task offloading in the event of partial system degradation or network instability (Kaleem et al., 2019; Dogea et al., 2023; Callegaro et al., 2021; Pacheco et al., 2021; Petrenko et al., 2023; El Sayed et al., 2023; Sara et al., 2016). Additionally, hybrid cloud-edge frameworks offer parallel computational paths to guarantee task execution continuity under localized failures or load fluctuations (Luo et al., 2020; La Salandra et al., 2024; Khan et al., 2022; Chen and Ma, 2024; Wang et al., 2020a; Shen et al., 2023a; Tang et al., 2023). These fault-tolerant architectures directly contribute to dependability attributes such as availability, reliability, and resilience by enabling uninterrupted operations and graceful degradation in the presence of faults, rather than complete system failure.
@@ 6.2.4. Network Topology Design
Network Topology Design refers to fault tolerance strategies that focus on maintaining reliable and resilient communication within systems, especially in environments where signal disruption, disconnection, or routing instability is common. Rather than preventing communication faults outright, these techniques aim to sustain service availability and network functionality even when individual links or nodes fail during operation. In UAV-based systems, dynamic topologies such as flying ad hoc networks (FANETs), mesh structures, and mobile relays are often employed to mitigate link degradation and maintain multi-hop connectivity. For instance, adaptive topology control and link prediction mechanisms are used to anticipate and respond to potential disconnections due to UAV mobility or energy depletion (Xu et al., 2021; Guruprasad et al., 2024; Zhang et al., 2023b; Vamvakas et al., 2019; Chen et al., 2022b; Wang et al., 2021c; Dao et al., 2021; Ozger et al., 2018; Su, 2021; Wang et al., 2022c).
UAV-aided reconfiguration is another popular approach, where drones are repositioned to act as temporary relays or mobile base stations to bridge coverage gaps and reduce the risk of total communication failure (Zhou et al., 2024b; Kong et al., 2022a; Zhang et al., 2021a; Mukherjee et al., 2020; Nwadiugwu et al., 2021; Mukherjee et al., 2023; Yue et al., 2018; Wang et al., 2020b; Zhou et al., 2024b; Wang et al., 2019a). These topology-aware fault tolerance mechanisms directly contribute to availability, connectivity, and system resilience, ensuring that UAV missions such as remote surveillance, disaster response, or public safety operations can continue despite disruptions in the communication infrastructure.
@@ 6.3. Fault Removal
| Diagnosis | (Xu and Zhai, 2020), (Venkatagiri et al., 2018), (Deng et al., 2021), (Jo et al., 2018), (Amin and Obermaisser, 2018), (Ramasamy et al., 2024), (Zhang and Zhang, 2024), (Wang et al., 2021b), (Wang et al., 2022a), (Wang et al., 2024c), (Kostoglotov et al., 2024), (Bobronnikov et al., 2022), (Bouzidi et al., 2020), (Amin and Obermaisser, 2018) | |
| Correction | (Gu et al., 2022), (Luo et al., 2022), (Bobronnikov et al., 2022) | |
@@ 6.3.2. Diagnosis:
Diagnosis primarily addresses software faults, hardware faults, and network disconnection, where faults may manifest as unexpected delays, abnormal UAV motion, or degraded communication quality. Several studies utilize behavioral and model-based diagnosis to identify operator-induced errors by comparing them with expected control models during simulated missions, providing insights into human-machine interaction errors and their safety implications (Xu and Zhai, 2020; Amin and Obermaisser, 2018). Secondly, cloud and edge-enabled anomaly detection frameworks are increasingly employed to enhance scalability and responsiveness, particularly for large-scale UAV applications such as power line inspection and predictive maintenance. These include distributed isolation forest algorithms and embedded monitoring tools that enable timely fault isolation while minimizing computational burden (Ramasamy et al., 2024; Zhang and Zhang, 2024; Wang et al., 2022a).
On the other hand, intelligent agents and software-based diagnostic systems support modular and autonomous operation by integrating distributed reasoning or software approximation techniques. For example, agent-based building inspection platforms and approximation-aware video summarization systems leverage real-time assessment capabilities to maintain operational resilience and energy efficiency (Venkatagiri et al., 2018; Jo et al., 2018). Meanwhile, cooperative and network-level diagnosis approaches utilize drone-assisted sensing and edge intelligence to detect anomalies within broader communication networks, including space-air-ground infrastructures, thereby ensuring systemic fault awareness and coordination (Deng et al., 2021; Wang et al., 2021b).
@@ 6.4. Fault Forecasting
| Fault Prevention | Ordinal Evaluation | (Bouzidi et al., 2020), (Guo et al., 2022b), (Hossain et al., 2017), (Manzini et al., 2023), (Fesenko et al., 2023), (Venkatagiri et al., 2018), (Kramer and Stubberud, 2021), (Cao et al., 2021) |
| Probabilistic Evaluation | (Sabino et al., 2024), (Ngoenriang et al., 2022). (Goel et al., 2021), (Tahat et al., 2020), (Ji et al., 2024), (Apostolopoulos et al., 2023), (Bai et al., 2023a), (Brito et al., 2021), (Falcão et al., 2024), (Machida and Andrade, 2021), (Kim and Kim, 2021), (Manzoor et al., 2021), (Thapliyal et al., 2024), (Machida et al., 2023), (Dai et al., 2018), (She and Ouyang, 2021), (Machida et al., 2023), (Chen et al., 2021a), (Yang et al., 2017), (Falcão et al., 2023), (Watanabe and Machida, 2022), (Schäfer et al., 2018), (Jeong et al., 2020), (Picano and Fantacci, 2024), (Kabashkin, 2023), (Yao and Ansari, 2018), (She et al., 2018), (Silva et al., 2024), (Hu and Seiler, 2015), (Falcão et al., 2023), (Li et al., 2019), (Bai et al., 2023a), (Venkatagiri et al., 2018), (Zhang, 2020) |
@@ 6.4.2. Probabilistic Evaluation:
In the context of dependability of UAV-based systems, probabilistic evaluation directly supports reliability analysis, availability estimation, and maintenance planning, offering a mathematical foundation for system robustness under varying workloads and fault scenarios. For example, studies that apply probabilistic models to assess component failure rates and communication link interruptions under uncertain conditions (Sabino et al., 2024; Ngoenriang et al., 2022; Tahat et al., 2020). Other studies integrate Markov models and stochastic processes to estimate system uptime and mission success rates (Ji et al., 2024; Apostolopoulos et al., 2023; Bai et al., 2023a). Bayesian networks and Monte Carlo simulations are used to analyze the effects of uncertain environmental conditions or resource variability on UAV operations (Falcão et al., 2024; Kim and Kim, 2021; Manzoor et al., 2021).
@@ 7.2. Constellation of satellites
To address these, reliability considerations involve energy-efficient designs, such as low-power transceivers and burst transmissions, as well as hybrid communication strategies that combine satellite and terrestrial links (Wang et al., 2019c) with fallback mechanisms. Robust signal resilience could be realised through adaptive modulation, latency-aware decision-making, and security measures. While satellite-enabled UAVs offer unprecedented operational range and flexibility, their deployment requires careful architectural design and robust fault-tolerance mechanisms to maintain dependable performance in dynamic and contested environments.
@@ 7.3. B5G/6G network integration
Future research should explore the design of UAV-based systems that are tailored to B5G/6G environments (Shrestha et al., 2021). In addition, joint optimization of communication and computing, including seamless handovers and UAV trajectory planning with connectivity guarantees, will be crucial to maintain QoS under highly dynamic aerial conditions. Integration with terahertz communication, intelligent reflecting surfaces (IRS), and network function virtualization (NFV) can further extend the coverage, reduce energy consumption, and improve resilience of UAV networks (Chen et al., 2021b).
@@ Acknowledgements.
- Akbari et al. (2024) Mohammad Akbari, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2024. AoI-aware energy-efficient SFC in UAV-aided smart agriculture using asynchronous federated learning. IEEE Open J. Commun. Soc. 5 (2024), 1222–1242.
@@
- Amin and Obermaisser (2018) Sarah Amin and Roman Obermaisser. 2018. A time-triggered scheduling algorithm for active diagnosis in heterogeneous distributed systems. In 2018 IEEE International Conference on Computational Science and Engineering (CSE) (Bucharest). IEEE.
@@
- Awada et al. (2021) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2021. Air-to-air collaborative learning: A multi-task orchestration in federated aerial computing. In 2021 IEEE 14th International Conference on Cloud Computing (CLOUD) (Chicago, IL, USA). IEEE.
- Awada et al. (2022) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2022. AirEdge: A dependency-aware multi-task orchestration in federated aerial computing. IEEE Trans. Veh. Technol. 71, 1 (Jan. 2022), 805–819.
- Awada et al. (2023) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, Shuangzhi Li, and Shouyi Yang. 2023. EdgeDrones: Co-scheduling of drones for multi-location aerial computing missions. J. Netw. Comput. Appl. 215, 103632 (June 2023), 103632.
@@
- Bai et al. (2023b) Haohao Bai, Zhineng Luo, Meihui Chen, and Senke Gui. 2023b. Energy saving research on auxiliary task unloading and resource allocation of UAV based on edge computing. In 2023 International Conference on Communications, Computing and Artificial Intelligence (CCCAI) (Shanghai, China). IEEE, 103–108.
@@
- Bakhtiari et al. (2020) Zohreh Bakhtiari, Rozita Jamili Oskouei, Mona Soleymani, and Akhtar Hussain Jalbani. 2020. Presenting an effective method to detect and track the broken path in VANET using UAVs. Wirel. Commun. Mob. Comput. 2020 (Aug. 2020), 1–12.
@@
- Chen et al. (2022b) Lunyuan Chen, Rui Zhao, Ke He, Zichao Zhao, and Liseng Fan. 2022b. Intelligent ubiquitous computing for future UAV-enabled MEC network systems. Cluster Comput. 25, 4 (Aug. 2022), 2417–2427.
@@
- Chen et al. (2021b) Zhi Chen, Boyu Ning, Chong Han, Zhongbao Tian, and Shaoqian Li. 2021b. Intelligent Reflecting Surface Assisted Terahertz Communications toward 6G. arXiv:2104.02897 [cs.IT] https://arxiv.org/abs/2104.02897
@@
- Chu et al. (2021) Liju Chu, Xuejun Li, Jia Xu, Azadeh Ghari Neiat, and Xiao Liu. 2021. A holistic service provision strategy for drone-as-a-service in MEC-based UAV delivery. In 2021 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE.
@@
- Falcão et al. (2023) Marcos Falcão, Caio Souza, Andson Balieiro, and Kelvin Dias. 2023. Dynamic resource allocation for URLLC in UAV-enabled multi-access edge computing. In 2023 Joint European Conference on Networks and Communications & 6G Summit (EuCNC/6G Summit) (Gothenburg, Sweden). IEEE.
@@
- Faraci et al. (2023) Giuseppe Faraci, Santi Agatino Rizzo, and Giovanni Schembra. 2023. Green edge intelligence for smart management of a FANET in disaster-recovery scenarios. IEEE Trans. Veh. Technol. 72, 3 (March 2023), 3819–3831.
@@
- Fatima et al. (2024) Nida Fatima, Paresh Saxena, and Giovanni Giambene. 2024. Deep reinforcement learning based computation offloading for xURLLC services with UAV-assisted IoT-based multi-access edge computing system. Wirel. Netw. 30, 9 (Dec. 2024), 7275–7291.
@@
- Gao et al. (2019) Ang Gao, Yansu Hu, Wei Liang, Yizhi Lin, Lixin Li, and Xu Li. 2019. A QoE-oriented scheduling scheme for energy-efficient computation offloading in UAV cloud system. IEEE Access 7 (2019), 68656–68668.
- Gao et al. (2021) Ang Gao, Qi Wang, Wei Liang, and Zhiguo Ding. 2021. Game combined multi-agent reinforcement learning approach for UAV assisted offloading. IEEE Trans. Veh. Technol. 70, 12 (Dec. 2021), 12888–12901.
@@
- Geng et al. (2024) Liwei Geng, Hongbo Zhao, and Changming Zou. 2024. Stackelberg game-based computation offloading and pricing in UAV assisted vehicular networks. IEEE Trans. Reliab. (2024), 1–15.
@@
- Gu et al. (2022) Xiaohui Gu, Guoan Zhang, Mingxing Wang, Wei Duan, Miaowen Wen, and Pin-Han Ho. 2022. UAV-aided energy-efficient edge computing networks: Security offloading optimization. IEEE Internet Things J. 9, 6 (March 2022), 4245–4258.